OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –5? y = –3x + 5 y = –3x – 5 y = 3x + 5 y = 3x – 5

OpenStudy (anonymous):

slope of -3 and a y-intercept of -5 y= -3x+5 y= -3x-5 y=3x+5 y=3x-5

OpenStudy (nurali):

y=mx+b m=slope b=y-intercept put in equation y=(-3)x+(-5) y=-3x-5

OpenStudy (anonymous):

thanks :) @Nurali

OpenStudy (nurali):

My Pleasure.